OpenStudy (anonymous):

Simplify the expression. 3(x + 3) – 2x + 4 A. 5x + 5 B. x + 13 C. 6x – 2 D. 3x – 11 I think its B but I just want to be sure

OpenStudy (anonymous):

3(x + 3) – 2x + 4 = 3x+3x3 - 2x +4 = x+13

OpenStudy (anonymous):

yes x+13

OpenStudy (anonymous):

kay thank you!!!